OpenStudy (anonymous):

Find the solutions to the equation (check all that apply) x2+14x+48=0. a.x=4 b.x=-8 c.x=6 d.x=-2 e.x=3 f.x=-6

OpenStudy (anonymous):

You can either plug in each of the values (a-f) in as x and see if you get 0 on the left...

OpenStudy (anonymous):

...or you can factor x^2+14x+48 and see what you get.

OpenStudy (anonymous):

i'd suggest factoring. do you know how to do that?

OpenStudy (anonymous):

like foiling?

OpenStudy (anonymous):

could you just help

OpenStudy (anonymous):

it's FOIL in reverse.

OpenStudy (anonymous):

your factored answer is going to look like (x + a)(x + b)... you just need to find the numbers a and b. if you foil (x +a)(x + b) out, you get x^2 +(a + b)x + ab...

OpenStudy (anonymous):

so to match what you have, you need two numbers that multiply to 48 but add to 14.

OpenStudy (anonymous):

I know and I did it and i cant find a solution i just plugged everything in and i still dont have a solution

OpenStudy (anonymous):

what two numbers multiply to 48 but add to 14?

OpenStudy (anonymous):

8 & 6\

OpenStudy (anonymous):

good. so a and b are 8 and 6. so you have (x + 8)(x + 6) = 0 as your factored form.

OpenStudy (anonymous):

now, if a * b = 0, then either a=0 or b=0.

OpenStudy (anonymous):

you have (x + 8) * (x + 6) = 0 so, either x+8 = 0 or x+6 = 0. Solve each of these for x and you get your two answers.

OpenStudy (anonymous):

x^2+14x+48 = 0 x^2+6x+8x+48 = 0 [48=6*8 and 6+8 =14(middle term)] x(x+6)+8(x+6)=0 [taking x common] (x+6)*(x+8)=0 now x=-6 or x=-8; ANS Both the values satisfy equation x^2+14x+48=0.

OpenStudy (anonymous):

Thank you!